OpenStudy (anonymous):

Find the value of y for the equation Y=-3x + 5, when x=-2

OpenStudy (callisto):

Put x=-2 into the equation. Y=-3(-2) + 5 = 6+5 = ?

OpenStudy (anonymous):

just plug it in.. Y=(-3)(-2)+5 =6+5 =11

OpenStudy (anonymous):

Thank You